OpenStudy (boynet):

Find the sum of the positive square free divisors of 5400. Note: A natural number is square free if it is not the multiple of any perfect square greater than 1.

OpenStudy (sshayer):

 now try to write required divisors

OpenStudy (sshayer):

i think only divisors are 2,3,5,6 which are not multiples of square numbers.

OpenStudy (sshayer):

10 is also there.

ganeshie8 (ganeshie8):

How about 15

OpenStudy (sshayer):

yes it is also there. I am thinking about more if any.

OpenStudy (sshayer):

30 is also there.

OpenStudy (sshayer):

so for we are able to find 2,3,5,6,10,15,30

OpenStudy (boynet):

I think I got it now. I figured out a formula. The prime factorization consists of 2,3, and 5. (2+1) * (3+1) * (5+1) = 3*4*6=72. I think thats the answer? Thanks for the help!

ganeshie8 (ganeshie8):

Nice, it seems we have to find the sum of all the divisors of \(2\times 3\times 5\)

ganeshie8 (ganeshie8):

72 looks good !

OpenStudy (boynet):

Thanks again